# High-energy physics image classification: A survey of Jet applications

### v1

@@ 3.1 Available datasets
The conseil Européen pour la recherche nucléaire (CERN) open data portal provides access to a variety of datasets from experiments conducted at the Large LHC. These datasets include information about collisions, particles, and Jet images. The portal offers a great starting point for those interested in HEP datasets. Figure 3 illustrates samples of Jet images, featuring the average of $p_{T}$-normalized quark and gluon Jet images across 5 distinct $\chi$ bins. The Jet images may undergo different preprocessing techniques, discussed later, prior to input into ML/DL models for classification or prediction tasks. Table 3 showcases the datasets most commonly utilized in the reviewed research within this paper, including some generated datasets.generated.
@@
Dataset/EG
Name
Description
DLA ?
Dataset
ATLAS open data
@@
Is one of the largest particle physics experiments at the LHC. They offer an "Open Data" initiative with datasets that include collision data and simulated samples. These datasets can be used to study Jet images and other particle physics phenomena.
Yes [^1] http://opendata.cern.ch/search?page=1&size=20&experiment=ATLAS
Dataset
CMS open data
@@
Compact muon solenoid (CMS) is another major experiment at the LHC. Similar to ATLAS, CMS provides open data for educational and research purposes. The datasets include information about collisions, particles, and Jets.
Yes  [^2] https://opendata.cern.ch/search?page=1&size=20&q=jet%20images&experiment=CMS
EG
Delphes simulation
@@
Is a particle physics event generator designed to produce simulated collision events that are similar to those observed in real experiments. It includes tools to generate Jet images based on the data produced in simulations.
Yes [^3] https://cp3.irmp.ucl.ac.be/projects/delphes
EG
MadGraph
@@
Is a popular event generator used in particle physics simulations. It can generate events involving Jets and other particles, which can then be turned into Jet images.
Yes [^4] http://madgraph.phys.ucl.ac.be/
EG
FASTSim
@@
Is a tool for simulating high-energy particle collisions. It can generate Jet images from simulated collision events and is often used for studying ML techniques in HEP.
Yes [^5] https://twiki.cern.ch/twiki/bin/view/CMSPublic/SWGuideFastSimulation
Dataset
Complete [8]
@@
It belongs to CERN and contains muon, kaon, pion, electron, and proton. In the complete dataset training, 400,000 Jets are used for training, and the remaining 290,000 are used for testing and assessing performance.
No
Dataset
Monte Carlo
@@
It is generated through a dependable framework, created by integrating various tools like Pythia 8 for generating high-energy physics events, Delphes for emulating the detector’s response, and RAVE for reconstructing secondary vertices. [12].
No
Dataset
Top tagging
@@
This dataset comprises 1.2 million training samples, 400,000 for validation, and another 400,000 for testing. Each entry in this dataset corresponds to an individual Jet, with its source being either an energetic top quark, a light quark, or a gluon. These events were generated using the PYTHIA8 Monte Carlo event generator, and the response of the ATLAS detector is simulated using the DELPHES software package
Yes [^6] https://zenodo.org/record/2603256
Dataset
Quark-gluon tagging
@@
The dataset is created by generating signal (quark) and background (gluon) Jets through PYTHIA8. For the signal Jets, the process involves $Z(\rightarrow νν)+(u,d,s)$, and for the background Jets, it uses $Z(\rightarrow νν)+g$. Notably, there is no simulation of the detector. The particles that are not neutrinos in the final state are grouped into Jets using the anti-kT algorithm with a radius parameter of $R=0.4$. In total, this dataset contains 2 million Jets, evenly split between signal and background categories [13].
No
Dataset
Higgs dataset
@@
The dataset originates from Monte Carlo simulations. The initial 21 attributes (found in columns 2-22) represent particle detector-derived kinematic properties within the accelerator. The remaining seven attributes are transformations of the initial 21, constituting high-level features engineered by physicists to aid in distinguishing between the two categories.
Yes [^7] https://archive.ics.uci.edu/dataset/280/higgs
Dataset
QCD multi-Jet
Certainly! Here’s the paraphrased version in LaTeX format:
[…]
No
@@ 3.2 Pre-processing of Jet images
The objective of preprocessing input data is to support the model in addressing an optimization challenge. Usually, these preprocessing actions are not mandatory, but they are employed to enhance the numerical convergence of the model, considering the real-world constraints imposed by limited datasets and model dimensions, along with the specific parameter initialization choices. In HEP, (i) $\eta$ represents pseudorapidity, which is a measure related to the polar angle of a particle’s trajectory. It is commonly used because it is less affected by relativistic effects and is approximately invariant under boosts along the beamline. (ii) $\phi$ represents the azimuthal angle, which is the angle around the beamline. (iii) Together, $\eta$ and $\phi$ provide a way to specify the direction and position of particles or energy deposits within the detector. These coordinates are particularly useful for representing and analyzing the distribution of particles produced in high-energy collisions. (iv) The combination of $\eta$ and $\phi$ can be thought of as a way to navigate and map the detector’s components in a way that is sensitive to the underlying physics processes. (v) $\eta-\phi$ space is a coordinate system used to describe the properties and positions of particles or objects within particle detectors, particularly in experiments at large colliders like the LHC.
@@ 3.3 Feature extraction and selection
Di Luca et al. [12] presents an automated feature selection procedure for particle Jet classification in HEP experiments. The authors use ML boosted tree algorithms to rank the importance of observables and select the most important features associated with a particle Jet. They apply this method to the specific case of boosted Higgs boson decaying to two b-quarks ($H\rightarrow bb$) tagging and demonstrate the impact of feature selection on the performance of the classifier to distinguish these events amidst the substantial and unalterable background originating from quantum chromodynamics (QCD) multi-Jet production. They also train a fully connected neural network to tag the Jets and compare the results obtained using all the features or only those selected from the procedure which consists of two main steps: data preparation and feature ranking extraction. The authors discover that the azimuthal angles of the large-R Jet and the variable radius (VR)-track Jets appear towards the end of the feature ranking. At the top of the ranking, they find the $p_{T}$ of the two VR-track Jets, along with certain details regarding the secondary vertex, such as its mass, energy, and displacement. The study shows that selecting the highest-ranked features achieves performance nearly as effective as that of the full model, with only a slight deviation of a few percent. This approach can be expanded to accommodate the increased number of observable variables that upcoming collider experiments will gather from high $p_{T}$ particle Jets. The data for this research comes from proton-proton collision events featuring a boosted Higgs boson that decays into two $b$ quarks. In [16], solutions have been proposed for classifying events extracted from the 2014 Higgs ML Kaggle dataset [^8] https://www.kaggle.com/c/higgs-boson. The dataset includes a mix of low-level and high-level attributes: it contains 18 low-level features that include three-dimensional momenta ($p_{T}$, $\eta$, $\phi$), missing transverse momentum, and the total transverse momentum from all jets; additionally, there are 13 high-level features motivated by physics, covering invariant masses and angular separations among objects in the final state. Table 4 summarizes the features utilized, which hold potential for future application within the context of HEP. The authors’ aim to ensure that the suggested networks make effective use of low-level information; otherwise, there’s a risk of losing these features during selection. Their focus lies in determining the necessity of high-level features. The proposed DNN model effectively utilize the low-level information in the data and autonomously learn their own high-level representations. Boost-invariant polynomial (BIP) features are a type of mathematical representation used in HEP for analyzing particle collision data. They are constructed to be invariant under boosts, meaning they remain unchanged under transformations to different reference frames with different velocities. These features are designed to capture important characteristics of particle jets, such as their energy distribution and substructure, while ensuring consistency across various experimental conditions. BIP features are particularly useful for tasks like jet tagging and classification in HEP experiments, as employed in [17].
@@
Table 4: Possible combinations of Jet features to generate new high- and low-level features that could potentially improve ML classification for Jet HEP. The performance of employing these features are in [16].
Level
Suggested feature name
Description
Grouping
DER_mass_MMC
@@
The Higgs boson’s mass was estimated using a hypothesis-driven fitting method.
Higgs, Mass
DER_mass_transverse_met_lep
@@
Transverse mass associated with the lepton and $P_{\text{miss}}^{T}$
Higgs, Mass
DER_mass_vis
The mass invariant to both the lepton and the tau
Higgs, Mass
High-level features
DER_pt_h
@@
Transverse momenta of the combined vector of the lepton, tau, and $P_{\text{miss}}^{T}$
Higgs, 3-momenta
DER_deltaeta_jet_jet
@@
Absolute disparity in pseudorapidity between the leading and subleading jets (undefined for less than two jets)
Jet with angular
@@
DER_mass_jet_jet
The invariant mass of the primary and secondary jets (not applicable when there are fewer than two jets).
Jet, Mass
DER_prodeta_jet_jet
The multiplication of the pseudo rapidities for the foremost and next-to-foremost jets (inapplicable if fewer than two jets are present).
Jet, 3-momenta
DER_deltar_tau_lep
Distance between the lepton and the tau in the $\eta-\phi$ plane.
Final state, Angular
DER_pt_tot
The $p_{T}$ resulting from the vector addition of the $p_{T}$ of the lepton, tau, the primary and secondary jets (when applicable), and $P_{\text{miss}}^{T}$.
Final-state, Sum
DER_sum_pt
Total transverse momentum of the lepton, tau, and all jets
global event, Sum
DER_pt_ratio_lep_tau
Ratio of the transverse momenta of the lepton to that of the tau
Final state, 3-momenta
DER_met_phi_centrality
Centrality of the azimuthal angle of $P_{\text{miss}}^{T}$ relative to the lepton and the tau
Final state, Angular
DER_lep_eta_centrality
The centrality measure of the lepton’s pseud-orapidity in comparison to the primary and secondary jets (not applicable for fewer than two jets).
Jet, Angular
PRI_tau_[px/py/pz]
The 3-momenta of the tau expressed in Cartesian coordinates.
Final state, 3-momenta
Low-level features
PRI_lep_[px/py/pz]
The lepton’s 3-momenta represented in Cartesian coordinates.
Final state, 3-momenta
PRI_met_[px/py]
The constituent parts of the missing transverse momentum vector expressed in Cartesian coordinates.
Final state, 3-momenta
PRI_met
The magnitude of the missing transverse momentum vector represented in Cartesian coordinates.
Final state, 3-momenta
PRI_met_sumet
Total sum of transverse energy.
Final-state, Energy
PRI_jet_num
Count of jets present in the event.
Jet, Multiplicity
PRI_jet_leading_[px/py/pz]
The three-dimensional momenta of the primary jet expressed in Cartesian coordinates (not applicable if there are no jets present).
Jet, 3-momenta
PRI_jet_subleading_[px/py/pz]
The 3-momenta of the secondary jet represented in Cartesian coordinates (not defined if fewer than two jets are present).
Jet, 3-momenta
PRI_jet_all_pt
Total sum of the transverse momenta of all jets in Cartesian coordinates.
Jet, 3-momenta
@@ 4 Available AI models for HEP Jet images
Many DL architectures have been proposed in the SOTA of HEP domain to identify particles, Table LABEL:tab:archCmp summarizes and compares the most efficient DL models, used in HEP, based on their architectures and performances. For example, the work in [18] proposed PartT, which is a new Transformer-based architecture for Jet tagging. Its main task is to identify the origin of a Jet of particles produced in HEP experiments. ParT makes use of two sets of inputs: (i) the particle input, which includes a list of features for every particle and forms an array, and (ii) the interaction input, which is a matrix of features for every pair of particles. ParT employs a novel pairwise multi-head attention (P-MHA) mechanism, which allows the model to attend to pairs of particles and learn their interactions. The P-MHA is more effective than standard plain multi-head attention. This assertion is substantiated when the pre-trained ParT models are fine-tuned on two widely adopted Jet tagging benchmarks, the quark-gluon tagging dataset and the binary classification dataset for identifying boosted $W$ bosons decaying to two quarks. The fine-tuning process involves training the ParT models on a smaller labeled dataset specific to each benchmark, which allows the models to learn the specific features and patterns relevant to each task. The fine-tuned ParT models achieve significantly higher tagging performance than the models trained from scratch and outperform the previous SOTA models, including ParticleNet and other Transformer-based models.
@@
Unlike the previous models, the authors in [19] proposed the equivariant graph neural networks (EGNN) model, which is a graph neural network (GNN) architecture that is translation, rotation, and reflection equivariant (E(n)), and permutation equivariant with respect to an input set of points. It uses a set of filters that are equivariant to the action of the symmetry group, which are constructed using a combination of radial basis functions and Chebyshev polynomials. The EGNN can retain the flexibility of GNNs while remaining E(n) equivariant as the radial field algorithm and without the need to compute expensive operations (i.e. spherical harmonics). The EGNN outperforms other equivariant and non-equivariant alternatives while still being efficient in terms of running time. Besides, EGNN reduces the error with respect to SOTA method by 32%.
@@
Another DL architecture called the DeepJet model, designed to consider the full information of all Jet constituents, including secondary vertices, global event variables, and charged and neutral particles, simultaneously, is proposed in [22]. The model architecture consists of several components: (i) Automatic feature extraction is conducted for each constituent through convolutional branches that include $1\times 1$ convolutional layers. Distinct convolutional branches are allocated for vertices, charged particle flow candidates, and neutral particle flow candidates. (ii) The output of the convolutional branches is used to construct a graph representation of the Jet, where each constituent is represented as a node in the graph. The edges between the nodes are determined by a distance metric that takes into account the kinematic properties of the constituents. (iii) The graph representation of the Jet is then processed by several graph convolutional layers, which are designed to capture the correlations between the constituents. The graph convolutional layers use a learnable filter that is applied to the graph representation of the Jet. (iv) The output of the graph convolutional layers is then fed into several dense layers, which are designed to perform the final classification task. The dense layers use a combination of fully connected and batch normalization layers. In the context of the DeepJet model, the recurrent neural network (RNN) layer is an important component of the DeepJet model, as it allows the model to capture the sequential information in the charged particle tracks and to use this information to improve the classification performance. The DeepJet model has been shown to achieve SOTA performance in Jet flavour classification and quark/gluon discrimination tasks. The model was tested using CMS simulation and was found to outperform previous classifiers, including the IP3D algorithm. The DeepJet model underwent a comparative analysis against a binary quark/gluon classifier from the CMS reconstruction framework. An improvement in performance was noted with the use of the DeepJet model on a dataset comprised exclusively of light quark and gluon Jets. Moreover, the DeepJet model was found to be more robust to variations in the Jet constituents and kinematics, which makes it more suitable for use in real-world scenarios. In terms of DeepJet’s Performance, using the function of reconstructed vertices, b-Jet efficiency can reach 92%, and when the function of Jet $p_{T}$, b-Jet efficiency is around 95%.
@@
Particle flow network with ID (PFN-ID) model [23] is another proposed type of ML architecture that takes particles as input and processes them in a way that is dependent on the order the particles were fed into the network. The PFN-ID architecture is based on the Deep Sets framework and includes full particle ID information. The Deep Sets framework is a ML approach that allows for learning directly from sets of features or "point clouds". The following are the main steps of the framework: (i) Map each element of the set to a latent space using a shared function. (ii) Aggregate the latent representations of the elements using a symmetric function. (iii) Map the aggregated latent representation back to the output space using a shared function. The framework guarantees that a generic symmetric function can be represented by an additive latent space. In the context of particle-level collider observables, each particle is mapped to a latent representation and then summed over, and observables are then functions on that latent space. This decomposition encompasses a wide variety of existing event- and Jet-level collider observables and representations, including image-based and moment-based methods. The PFN-ID improves the classification performance of the particle flow network (PFN) model for discriminating quark and gluon Jets. Results show that PFN-ID slightly outperforms RNN-ID, whereas the PFN and RNN are comparable.
@@
Figure 12: The architecture of TopoDNN model suggested in [25].
Ref.
Year
Model
Performance TT Performance QG Other performance
Link
Acc.
AUC
Acc
AUC
Acc
AUC
[18]
2022
partT
0.944
0.9877
0.852
0.9230
Yes [^9] https://github.com/jet-universe/particle_transformer
[13]
2022
LorentzNet
0.942
0.9868
0.844
0.9156
No
[19]
2021
EGNN
0.922
0.9760
0.803
0.8806
Yes [^10] https://github.com/vgsatorras/egnn
[20]
2020
ParticleNet
0.940
0.9858
0.840
0.9116
No
[21]
2020
LGN
0.929
0.9640
0.803
0.8324
Yes [^11] https://github.com/fizisist/LorentzGroupNetwork
[23]
2019
PFN-ID
0.932
0.9819
0.9005
No
[24]
2018
CNN tagger
0.87 (DTJ)
0.943 (DTJ)
No
0.945 (CJ)
0.988 (CJ)
[25]
2017
TopoDNN
0.916
0.972
No

### v2

@@ 3.1 Available datasets
The conseil Européen pour la recherche nucléaire (CERN) open data portal provides access to a variety of datasets from experiments conducted at the Large LHC. These datasets include information about collisions, particles, and Jet images. The portal offers a great starting point for those interested in HEP datasets. Figure 3 illustrates samples of Jet images, featuring the average of $p_{T}$-normalized quark and gluon Jet images across 5 distinct $\chi$ bins. The Jet images may undergo different preprocessing techniques, discussed later, prior to input into ML/DL models for classification or prediction tasks. Table LABEL:tab:datasets showcases the datasets most commonly utilized in the reviewed research within this paper, including some generated datasets.generated.
@@
Dataset/EG Name Description DLA ?
@@
Dataset ATLAS open data Is one of the largest particle physics experiments at the LHC. They offer an "Open Data" initiative with datasets that include collision data and simulated samples. These datasets can be used to study Jet images and other particle physics phenomena. Yes [^1] http://opendata.cern.ch/search?page=1&size=20&experiment=ATLAS
@@
Dataset CMS open data Compact muon solenoid (CMS) is another major experiment at the LHC. Similar to ATLAS, CMS provides open data for educational and research purposes. The datasets include information about collisions, particles, and Jets. Yes  [^2] https://opendata.cern.ch/search?page=1&size=20&q=jet%20images&experiment=CMS
@@
EG Delphes simulation Is a particle physics event generator designed to produce simulated collision events that are similar to those observed in real experiments. It includes tools to generate Jet images based on the data produced in simulations. Yes [^3] https://cp3.irmp.ucl.ac.be/projects/delphes
@@
EG MadGraph Is a popular event generator used in particle physics simulations. It can generate events involving Jets and other particles, which can then be turned into Jet images. Yes [^4] http://madgraph.phys.ucl.ac.be/
@@
EG FASTSim Is a tool for simulating high-energy particle collisions. It can generate Jet images from simulated collision events and is often used for studying ML techniques in HEP. Yes [^5] https://twiki.cern.ch/twiki/bin/view/CMSPublic/SWGuideFastSimulation
@@
Dataset Complete It belongs to CERN and contains muon, kaon, pion, electron, and proton. In the complete dataset training, 400,000 Jets are used for training, and the remaining 290,000 are used for testing and assessing performance [25]. No
@@
Dataset Monte Carlo It is generated through a dependable framework, created by integrating various tools like Pythia 8 for generating HEP events, Delphes for emulating the detector’s response, and RAVE for reconstructing secondary vertices [29]. No
@@
Dataset Top tagging This dataset comprises 1.2 million training samples, 400,000 for validation, and another 400,000 for testing. Each entry in this dataset corresponds to an individual Jet, with its source being either an energetic top quark, a light quark, or a gluon. These events were generated using the PYTHIA8 Monte Carlo event generator, and the response of the ATLAS detector is simulated using the DELPHES software package Yes [^6] https://zenodo.org/record/2603256
@@
Dataset Quark-gluon tagging The dataset is created by generating signal (quark) and background (gluon) Jets through PYTHIA8. For the signal Jets, the process involves $Z(\rightarrow νν)+(u,d,s)$, and for the background Jets, it uses $Z(\rightarrow νν)+g$. Notably, there is no simulation of the detector. The particles that are not neutrinos in the final state are grouped into Jets using the anti-kT algorithm with a radius parameter of $R=0.4$. In total, this dataset contains 2 million Jets, evenly split between signal and background categories [30]. No
@@
Dataset Higgs dataset The dataset originates from Monte Carlo simulations. The initial 21 attributes (found in columns 2-22) represent particle detector-derived kinematic properties within the accelerator. The remaining seven attributes are transformations of the initial 21, constituting high-level features engineered by physicists to aid in distinguishing between the two categories. Yes [^7] https://archive.ics.uci.edu/dataset/280/higgs
Dataset QCD multi-Jet Samples are generated across different ranges of scalar sum of $p_{T}$, namely 1000-1500 GeV, 1500-2000 GeV, and 2000-Inf GeV. After excluding samples with $p_{T}$ values less than 1000 GeV, the dataset consists of around 450. 103 training images, 150.103 validation images, and 150.103 testing images [16]. No
@@ 3.2 Pre-processing of Jet images
The objective of preprocessing input data is to support the model in addressing an optimization challenge. Usually, these preprocessing actions are not mandatory, but they are employed to enhance the numerical convergence of the model, considering the real-world constraints imposed by limited datasets and model dimensions, along with the specific parameter initialization choices. In HEP, (i) $\eta$ represents pseudorapidity, which is a measure related to the polar angle of a particle’s trajectory. It is commonly used because it is less affected by relativistic effects and is approximately invariant under boosts along the beamline, (ii) $\phi$ represents the azimuthal angle, which is the angle around the beamline, (iii) together, $\eta$ and $\phi$ provide a way to specify the direction and position of particles or energy deposits within the detector. These coordinates are particularly useful for representing and analyzing the distribution of particles produced in high-energy collisions, (iv) the combination of $\eta$ and $\phi$ can be thought of as a way to navigate and map the detector’s components in a way that is sensitive to the underlying physics processes, (v) $\eta-\phi$ space is a coordinate system used to describe the properties and positions of particles or objects within particle detectors, particularly in experiments at large colliders like the LHC.
@@ 3.3 Feature extraction and selection
Di Luca et al. [29] presents an automated feature selection procedure for particle Jet classification in HEP experiments. The authors use ML boosted tree algorithms to rank the importance of observables and select the most important features associated with a particle Jet. They apply this method to the specific case of boosted Higgs boson decaying to two b-quarks ($H\rightarrow bb$) tagging and demonstrate the impact of feature selection on the performance of the classifier to distinguish these events amidst the substantial and unalterable background originating from quantum chromodynamics (QCD) multi-Jet production. They also train a fully connected neural network to tag the Jets and compare the results obtained using all the features or only those selected from the procedure which consists of two main steps: data preparation and feature ranking extraction. The authors discover that the azimuthal angles of the large-R Jet and the variable radius (VR)-track Jets appear towards the end of the feature ranking. At the top of the ranking, they find the $p_{T}$ of the two VR-track Jets, along with certain details regarding the secondary vertex, such as its mass, energy, and displacement. The study shows that selecting the highest-ranked features achieves performance nearly as effective as that of the full model, with only a slight deviation of a few percent. This approach can be expanded to accommodate the increased number of observable variables that upcoming collider experiments will gather from high $p_{T}$ particle Jets. The data for this research comes from proton-proton collision events featuring a boosted Higgs boson that decays into two $b$ quarks. In [33], solutions have been proposed for classifying events extracted from the 2014 Higgs ML Kaggle dataset [^8] https://www.kaggle.com/c/higgs-boson. The dataset includes a mix of low-level and high-level attributes: it contains 18 low-level features that include three-dimensional momenta ($p_{T}$, $\eta$, $\phi$), missing transverse momentum, and the total transverse momentum from all jets; additionally, there are 13 high-level features motivated by physics, covering invariant masses and angular separations among objects in the final state. Table LABEL:tab:features summarizes the features utilized, which hold potential for future application within the context of HEP. The authors aim to ensure that the suggested networks make effective use of low-level information; otherwise, there’s a risk of losing these features during selection. Their focus lies in determining the necessity of high-level features. The proposed DNN model effectively utilize the low-level information in the data and autonomously learn their own high-level representations. Boost-invariant polynomial (BIP) features are a type of mathematical representation used in HEP for analyzing particle collision data. They are constructed to be invariant under boosts, meaning they remain unchanged under transformations to different reference frames with different velocities. These features are designed to capture important characteristics of particle jets, such as their energy distribution and substructure, while ensuring consistency across various experimental conditions. BIP features are particularly useful for tasks like jet tagging and classification in HEP experiments, as employed in [34].
@@
Table 4: Possible combinations of Jet features to generate new high- and low-level features that could potentially improve ML classification for Jet HEP. The performance of employing these features are presented in [33].
Level Suggested feature name Description Grouping
@@
DER_mass_MMC The Higgs boson’s mass was estimated using a hypothesis-driven fitting method Higgs, Mass
@@
DER_mass_transverse_met_lep Transverse mass associated with the lepton and $P_{\text{miss}}^{T}$ Higgs, Mass
DER_mass_vis The mass invariant to both the lepton and the tau Higgs, Mass
@@
High-level features DER_pt_h Transverse momenta of the combined vector of the lepton, tau, and $P_{\text{miss}}^{T}$ Higgs, 3-momenta
@@
DER_deltaeta_jet_jet Absolute disparity in pseudorapidity between the leading and subleading jets (undefined for less than two jets) Jet with angular
@@
DER_mass_jet_jet The invariant mass of the primary and secondary jets (not applicable when there are fewer than two jets) Jet, Mass
DER_prodeta_jet_jet The multiplication of the pseudo rapidities for the foremost and next-to-foremost jets (inapplicable if fewer than two jets are present) Jet, 3-momenta
DER_deltar_tau_lep Distance between the lepton and the tau in the $\eta-\phi$ plane Final state, Angular
DER_pt_tot The $p_{T}$ resulting from the vector addition of the $p_{T}$ of the lepton, tau, the primary and secondary jets (when applicable), and $P_{\text{miss}}^{T}$ Final-state, Sum
DER_sum_pt Total transverse momentum of the lepton, tau, and all jets global event, Sum
DER_pt_ratio_lep_tau Ratio of the transverse momenta of the lepton to that of the tau Final state, 3-momenta
DER_met_phi_centrality Centrality of the azimuthal angle of $P_{\text{miss}}^{T}$ relative to the lepton and the tau Final state, Angular
DER_lep_eta_centrality The centrality measure of the lepton’s pseud-orapidity in comparison to the primary and secondary jets (not applicable for fewer than two jets) Jet, Angular
PRI_tau_[px/py/pz] The 3-momenta of the tau expressed in Cartesian coordinates Final state, 3-momenta
Low-level features PRI_lep_[px/py/pz] The lepton’s 3-momenta represented in Cartesian coordinates Final state, 3-momenta
PRI_met_[px/py] The constituent parts of the missing transverse momentum vector expressed in Cartesian coordinates Final state, 3-momenta
PRI_met The magnitude of the missing transverse momentum vector represented in Cartesian coordinates Final state, 3-momenta
PRI_met_sumet Total sum of transverse energy Final-state, Energy
PRI_jet_num Count of jets present in the event Jet, Multiplicity
PRI_jet_leading_[px/py/pz] The three-dimensional momenta of the primary jet expressed in Cartesian coordinates (not applicable if there are no jets present) Jet, 3-momenta
PRI_jet_subleading_[px/py/pz] The 3-momenta of the secondary jet represented in Cartesian coordinates (not defined if fewer than two jets are present) Jet, 3-momenta
PRI_jet_all_pt Total sum of the transverse momenta of all jets in Cartesian coordinates Jet, 3-momenta
@@ 4 Available AI models for HEP Jet images
Many DL architectures have been proposed in the SOTA of HEP domain to identify particles, Table 5 summarizes and compares the most efficient DL models, used in HEP, based on their architectures and performances. For example, the work in [35] proposed PartT, which is a new Transformer-based architecture for Jet tagging. Its main task is to identify the origin of a Jet of particles produced in HEP experiments. ParT makes use of two sets of inputs: (i) the particle input, which includes a list of features for every particle and forms an array, and (ii) the interaction input, which is a matrix of features for every pair of particles. ParT employs a novel pairwise multi-head attention (P-MHA) mechanism, which allows the model to attend to pairs of particles and learn their interactions. The P-MHA is more effective than standard plain multi-head attention. This assertion is substantiated when the pre-trained ParT models are fine-tuned on two widely adopted Jet tagging benchmarks, the quark-gluon tagging dataset and the binary classification dataset for identifying boosted $W$ bosons decaying to two quarks. The fine-tuning process involves training the ParT models on a smaller labeled dataset specific to each benchmark, which allows the models to learn the specific features and patterns relevant to each task. The fine-tuned ParT models achieve significantly higher tagging performance than the models trained from scratch and outperform the previous SOTA models, including ParticleNet and other Transformer-based models.
@@
Unlike the previous models, the authors in [36] proposed the equivariant graph neural networks (EGNN) model, which is a graph neural network (GNN) architecture that is translation, rotation, and reflection equivariant (E(n)), and permutation equivariant with respect to an input set of points. It uses a set of filters that are equivariant to the action of the symmetry group, which are constructed using a combination of radial basis functions and Chebyshev polynomials. The EGNN algorithm possesses the same flexibility as the GNN technique, while also maintaining E(n) equivariance similar to the radial field algorithm. Additionally, it eliminates the requirement for computationally intensive procedures, such as spherical harmonics. The EGNN exceeds other equivariant and non-equivariant options while maintaining efficiency in terms of running time. Moreover, the EGNN approach demonstrates a 32% reduction in error compared to the SOTA method.
@@
Another DL architecture called the DeepJet model, designed to consider the full information of all Jet constituents, including secondary vertices, global event variables, and charged and neutral particles, simultaneously, is proposed in [39]. The model architecture consists of several components: (i) Automatic feature extraction is conducted for each constituent through convolutional branches that include $1\times 1$ convolutional layers. Distinct convolutional branches are allocated for vertices, charged particle flow candidates, and neutral particle flow candidates, (ii) the output of the convolutional branches is used to construct a graph representation of the Jet, where each constituent is represented as a node in the graph. The edges between the nodes are determined by a distance metric that takes into account the kinematic properties of the constituents, (iii) the graph representation of the Jet is then processed by several graph convolutional layers, which are designed to capture the correlations between the constituents. The graph convolutional layers use a learnable filter that is applied to the graph representation of the Jet, and (iv) the output of the graph convolutional layers is then fed into several dense layers, which are designed to perform the final classification task. The dense layers use a combination of fully connected and batch normalization layers. In the context of the DeepJet model, the recurrent neural network (RNN) layer is an important component of the DeepJet model, as it allows the model to capture the sequential information in the charged particle tracks and to use this information to improve the classification performance. The DeepJet model has been shown to achieve SOTA performance in Jet flavour classification and quark/gluon discrimination tasks. The model was tested using CMS simulation and was found to outperform previous classifiers, including the IP3D algorithm. The DeepJet model underwent a comparative analysis against a binary quark/gluon classifier from the CMS reconstruction framework. An improvement in performance was noted with the use of the DeepJet model on a dataset comprised exclusively of light quark and gluon Jets. Moreover, the DeepJet model was found to be more robust to variations in the Jet constituents and kinematics, which makes it more suitable for use in real-world scenarios. In terms of DeepJet’s Performance, using the function of reconstructed vertices, b-Jet efficiency can reach 92%, and when the function of Jet $p_{T}$, b-Jet efficiency is around 95%.
@@
Particle flow network with ID (PFN-ID) model [40] is another proposed type of ML architecture that takes particles as input and processes them in a way that is dependent on the order the particles were fed into the network. The PFN-ID architecture is based on the Deep Sets framework and includes full particle ID information. The Deep Sets framework is a ML approach that allows for learning directly from sets of features or "point clouds". The following are the main steps of the framework: (i) Map each element of the set to a latent space using a shared function. (ii) Aggregate the latent representations of the elements using a symmetric function. (iii) Map the aggregated latent representation back to the output space using a shared function. An additive latent space can be used to express a general symmetric function, as provided by the framework. Within the scope of particle-level collider observables, the process involves mapping each particle to a latent representation, which is subsequently collected. Subsequently, the observables are expressed as functions on this latent space. This decomposition includes a diverse range of current collider observables and representations at the event and Jet levels, including as image-based and moment-based techniques. The PFN-ID improves the classification performance of the particle flow network (PFN) model for discriminating quark and gluon Jets. Results show that PFN-ID slightly outperforms RNN-ID, whereas the PFN and RNN are comparable.
@@
Figure 12: The architecture of the TopoDNN model, consists of 4 layers with 300, 102, 12, and 6 nodes, respectively [42].
Table 5: A summary of available DL architectures for Jet HEP classification.
Ref. Year Model Performance TT Performance QG Other performance Link
Acc. AUC Acc AUC Acc AUC
[35] 2022 partT 0.944 0.9877 0.852 0.9230 Yes [^9] https://github.com/jet-universe/particle_transformer
[30] 2022 LorentzNet 0.942 0.9868 0.844 0.9156 No
[36] 2021 EGNN 0.922 0.9760 0.803 0.8806 Yes [^10] https://github.com/vgsatorras/egnn
[37] 2020 ParticleNet 0.940 0.9858 0.840 0.9116 No
[38] 2020 LGN 0.929 0.9640 0.803 0.8324 Yes [^11] https://github.com/fizisist/LorentzGroupNetwork
[40] 2019 PFN-ID 0.932 0.9819 0.9005 No
[41] 2018 CNN tagger 0.87 (DTJ) 0.943 (DTJ) No
0.945 (CJ) 0.988 (CJ)
[42] 2017 TopoDNN 0.916 0.972 No